ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 125, 2001

Alestalo, P., D. A. Trotsenko and J. Väisälä Isometric approximation

61

Andretta, A., I. Neeman and J. Steel The domestic levels of K^{c} are iterable

157

Borrelli, V. New examples of Lagrangian rigidity  221

Dajczer, M., L. A. Florit and R. Tojeiro On a class of submanifolds carrying an extrinsic totally umbilical foliation

203

Davies, B. and A. Neeman Algebraic geometry of the three-state chiral Potts model

253

Florit, L. A. See M. Dajczer

Friedgut, E. and V. Rödl Proof of a hypercontractive estimate via entropy

369

Gonçalves, J. Z. and D. S. Passman Unitary units in group algebras

131

Guesmia, A. Exact controllability for the wave equation with variable coefficients

83

Hamkins, J. D. Gap forcing  237

Kamae, T. Stochastic analysis based on deterministic Brownian motion

317

Lev, V. F., T. Łuczak and T. Schoen Sum-free sets in abelian groups

347

Łuczak, T. See V. F. Lev

Mauldin, R. D. and M. Urbański Gibbs states on the symbolic space over an infinite alphabet

93

Mitra, S. Teichmüller contraction in the Teichmüller space of a closed set in the sphere

45

Neeman, A. See B. Davies

Neeman, I. See A. Andretta

Passman, D. S. See J. Z. Conçalves

Preiss, D. and L. Zajíček Directional derivatives of Lipschitz functions

1

Rödl, V. See E. Friedgut Scherback, I. and E. Shustin Equimultiple deformations of isolated

singularities

293

Schoen, T. See V. F. Lev Shalev, A. Random generation of finite simple groups by p-regular or

p-singular elements

53

Shustin, E. See I. Scherback Steel, J. See A. Andretta

Steif, J. E. The T, T^{−1}-process, finitary codings and weak Bernoulli  29

Tojeiro, R. See M. Dajczer Trotsenko, D. A. See P. Alestalo Urbański, M. See R. D. Mauldin Väisälä, J. See P. Alestalo Zajíček, L. See D. Preiss